- 2. 5 c) A light bulb draws a current of 12A when connected to a 24-V battery. Find the bulbs resistance.
Answer #1

Given

Current I =12A

Voltage v = 24 V

From the ohm's law

Resistance R= v/I= 24/12=2ohm